OpenStudy (anonymous):

Factor: (2x+1)(3x+2)+(3x+5)(x-2) This one is really long and it's getting confusing. :(

OpenStudy (anonymous):

start with (2x+1)(3x+2)

OpenStudy (anonymous):

That's the problem... where do you start?!?!

OpenStudy (anonymous):

I mean can you factor that out

OpenStudy (anonymous):

What is 2x times 3x

OpenStudy (anonymous):

Would it be 6x^2?

OpenStudy (anonymous):

correct. Excellent job. Now it goes to 2x times two.

OpenStudy (anonymous):

4x?

OpenStudy (anonymous):

Yes fantastic. After that it is 1 times 3x. Are you seeing the pattern

OpenStudy (anonymous):

I think so? So would it be 3x+2?

OpenStudy (anonymous):

All together would it be (6x^2+4x)(3x+2)? -For the left side?

OpenStudy (anonymous):

Yes other than the fact that you do not multiply them together. Since you are combining them together you add them. Very fine job

OpenStudy (mew55):

9x^2+6x-8

OpenStudy (anonymous):

So would my final answer be -9x^2+12?

OpenStudy (anonymous):

No. That up there is the correct final answer but the leftside will be. 6x^2+7x+2

OpenStudy (mew55):

no because u dont have ur x value as well. look at my post. its 9x^2+6x-8 and simplifying it will give u (3x-2)(3x+4)

OpenStudy (anonymous):

The right side will be 3x^2-x-10

OpenStudy (anonymous):

How did you simplify it down into that answer mew55?

OpenStudy (anonymous):

You combine the two polynomials

OpenStudy (mew55):

okay so u know x*x=x^2 and x*1=x from the 9x^2+6x-8 that is simplified to (3x-2)(3x+4) it is 9x^2+12x-6x-8 its distribution, which is multiplying each number on one by by the other

OpenStudy (anonymous):

6x^2+7x+2+3x^2-x-10=9x^2+7x-5